OpenStudy (anonymous):

Two angles in a triangle have radian measure π/2 and π/5. What is the radian measure of the third angle? (Express your answer in terms of π.)

OpenStudy (anonymous):

The sum of the three angles in any triangle is \(\pi\) radians (180 deg). So, subtract the two given angles from \(\pi\) and what's left is the measure of the third angle.

OpenStudy (anonymous):

would it be 3 π?

OpenStudy (anonymous):

I don't really understand

OpenStudy (anonymous):

Can't be. The sum of all three angles has to be \(\pi\).

OpenStudy (anonymous):

\[\pi - \frac{ \pi }{ 2 } - \frac{ \pi }{ 5 } = ?\]

OpenStudy (anonymous):

What would the top answer be?

OpenStudy (anonymous):

You tell me. Remember, in order to subtract fractions, they must have the same denominator. What's the LCD here?

OpenStudy (anonymous):

Would the answer be 3 π/10?

OpenStudy (anonymous):

Correct. Very good!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome.